Lura:

Evaluate the expression shown below and write your answer as a fraction in simplest form. \frac{1}{2}\div-\frac{6}{5} 2 1 ​ ÷− 5 6 ​

Lura:

1/2 divided by -6/5

jhonyy9:

use this \[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b }×\frac{ d }{ c }\]